OpenStudy (anonymous):

[2.05] In a certain county, the number of Charter Schools is six less than twice the number of Alternative Schools. We know that there are 52 Charter Schools in the county. How many Alternative Schools are in the county? 58 44 29 23

OpenStudy (whpalmer4):

A = alternative schools C = charter schools If charter schools are 6 less than twice the number of alternative schools, then we can write C = 2*A - 6 If we also know that C = 52, then 52 = 2A - 6 and you should be able to solve that for the value of A.

OpenStudy (anonymous):

52-6=46 46/2=23 a=23??

OpenStudy (whpalmer4):

Let's try that out. If there are 23 alternative schools, and the number of charter schools is 6 less than twice 23, does that equal 52? 2*23 = 46 - 6 = 40. Maybe not...

OpenStudy (whpalmer4):

\[52= 2A -6\]Add 6 to both sides of the equation \[52+6=2A - 6 + 6\]\[58=2A\]\[A=\]

OpenStudy (anonymous):

a=29?

OpenStudy (whpalmer4):

Don't ask me, try it out. Does it solve the problem, or not?